OpenStudy (anonymous):

Nagaraj has a party-length sandwich 63in. long. The middle piece will be 4 in. longer than the shortest piece and the shortest piece will be 8in. shorter than the longest piece. How long will the three pieces be?

OpenStudy (anonymous):

 Say, the sandwich is divided into 3 lengths A, B and C. Now the middle piece is 8in. longer than the shortest piece. Let the longest piece be A (doesn't matter if it's A or C since you can rotate the sandwich around and it'll be at the other end). The shortest piece is C. Using the information given, we can now make some equations: B=4+C C=A-8 Now together, A+B+C=63, so do you know what to do from there?